OpenStudy (anonymous):

Which correctly describes how the graph of the inequality -4y - x ≥ 7 is shaded? Choices: Above the solid line Below the solid line Above the dashed line Below the dashed line

OpenStudy (danjs):

first you should solve that for y

OpenStudy (danjs):

\[-4y \ge x+7\]

OpenStudy (danjs):

dividing by a negative 4 will flip the sign around

OpenStudy (danjs):

\[y \le \frac{ -1 }{ 4 }x - \frac{ 7 }{ 4 }\]

OpenStudy (danjs):

\[\le ~~ and \ge~~~solid lines \] < and > are dotted lines, not including the numbers ON the line

OpenStudy (danjs):

so it will be a solid line, shaded below, from the less than or equal to

OpenStudy (anonymous):

Okay thank you :)

OpenStudy (danjs):

do you understand all that?

OpenStudy (anonymous):

yes a little more than I did before